# COEF-VQ: Cost-Efficient Video Quality Understanding through a Cascaded Multimodal LLM Framework

### v1

@@ 3.3. Ablation Studies and Analysis
Comparisons between Audio Fusion Strategies. In this section, we compare the performance of Late fusion used in our framework, with Early fusion, as shown in Table 3. In the Early fusion approach, the output from the audio encoder is fed into the LLM as the final input token. The results indicate that Late fusion achieves a slightly higher F1 score. Our interpretation is that even though Early fusion would provide a stronger feature crossing, it requires a foundation model with better modality alignment between audio and VL, which currently there isn’t one. Therefore, we ended up integrating the audio feature with Late fusion due to its proven effectiveness, as well as robustness in terms of audio feature coverage.
@@ 4. Related Work
Additionally, SigLIP (Zhai et al., 2023) introduces a simplified approach by employing a pairwise sigmoid loss for image-text pre-training, providing an alternative to conventional contrastive objectives. BEiT3 (Wang et al., 2022) proposes a general-purpose multimodal foundation model that adopts a multi-way transformer architecture, integrating both masked image modeling and masked language modeling objectives to support a broad array of vision-language tasks. This model is designed to serve as a versatile framework, accommodating various types of multimodal data and allowing for both robust image understanding and sophisticated text comprehension.
Collectively, these models and approaches represent a new era of VLP, pushing the boundaries of vision-language understanding across a range of applications. By refining alignment mechanisms, introducing more efficient training objectives, and expanding the capability of multimodal transformers, these advancements pave the way for more comprehensive and flexible models capable of handling complex vision-language tasks at scale.
@@
To address this gap, LLaVA-OneVision(Li et al., 2024) has been introduced with the goal of achieving top-tier performance across single-image, multi-image, and video tasks, thereby exhibiting robust generalization and effective transfer of learned features in cross-scenario. By leveraging advanced techniques for cross-modal alignment and representation, LLaVA-OneVision enables flexible adaptation and accurate predictions across different visual contexts, such as switching from high-detail image analysis to complex, temporally sequenced video tasks.
Several other versatile open models have also shown promising potential for multi-scenario applications in this space. Video-LLaMA(Zhang et al., 2023) proposes a multi-modal framework that empowers Large Language Models (LLMs) with the capability of understanding both visual and auditory content in the video. VILA(Lin et al., 2024) examined effective pretraining strategies to adapt LLMs for vision-related tasks. InternLMX-Composer-2.5(Zhang et al., 2024) offers advanced capabilities for handling long-context input and output, supporting features like ultra-high resolution image comprehension, detailed video analysis, multi-turn multi-image dialogue, webpage creation, and article composition. Meanwhile, Qwen2-VL(Wang et al., 2024) introduces the Naive Dynamic Resolution mechanism, enabling the model to dynamically process images of varying resolutions into different quantities of visual tokens. This approach enhances the model’s ability to generate efficient and precise visual representations, closely mimicking human perception.
@@ 5. Conclusion
Modern video understanding models benefit significantly from a multimodal approach. By leveraging the complementary strengths of these modalities, multi-tower architectures achieve robust and holistic performance across general tasks. To better address the nuanced demands of specialized applications, a unified Multimodal LLM framework can further improve cross-modal interactions and sensitivity to domain-specific cues, such as assessing video quality or policy compliance on platforms like TikTok, surpassing traditional multi-tower limitations.
Given the substantial GPU demands of Multimodal LLMs, especially in real-time scenarios, we propose a cascade structure for efficient deployment. This two-stage model uses a lightweight classifier initially, filtering videos before engaging the MLLM for more in-depth consideration on selected cases, thus reducing resource usage while maintaining accuracy. Empirical results show that this cascade approach not only optimizes GPU use but also enhances classification precision for specialized tasks, marking a substantial advancement in scalable multimodal video classification.
@@ 6. Future Exploration
For future exploration, a promising direction is to expand the COEF-VQ framework into a more versatile, multi-issue MLLM that can handle diverse video quality issues within a single, foundational model. By building a comprehensive model capable of addressing multiple video assessment tasks (e.g., quality classification, content moderation, and genre classification), platforms like TikTok could achieve more streamlined processing with less redundancy across models. This unified approach could also improve consistency in video understanding and reduce the need for task-specific model deployment, optimizing both resource and maintenance efforts.
Another important step involves improving audio modality integration. Currently, the framework relies on late-fused feature concatenation, which may not fully capture the nuanced alignment between text, audio and visual cues. To address this, future versions of COEF-VQ could employ modality alignment techniques to establish stronger interactions between text, audio and visual information. This enhanced integration would allow the model to better interpret scenarios where audio cues are critical to understanding video quality, such as distinguishing between background noise and intentional soundtrack effects, ultimately leading to more accurate quality assessments.

### v2

@@ 3.3. Ablation Studies and Analysis
Comparisons between Audio Fusion Strategies. We compare the performance of Late fusion used in our framework, with Early fusion, as shown in Table 3. In the Early fusion approach, the output from the audio encoder is fed into the LLM as the final input token. The results indicate that Late fusion achieves a slightly higher F1 score. Our interpretation is that even though Early fusion would provide a stronger feature crossing, it requires a foundation model with better modality alignment between audio and VL, which currently there isn’t one. Therefore, we ended up integrating the audio feature with Late fusion due to its proven effectiveness, as well as robustness in terms of audio feature coverage.
Comparisons between Filtering Criteria. We evaluate the performance of the Entropy-based filtering approach in our framework by comparing it with the Confidence-based approach, as presented in Table 6. To ensure a fair comparison, we maintain a consistent QPS ratio for both methods. The Confidence-based approach selects samples with confidence scores exceeding a predefined threshold.
The results show that the Entropy-based filter achieves higher recall across different precision levels, indicating its effectiveness in selecting more informative cases for the second stage. This demonstrates that our Entropy-based method prioritizes uncertain cases more efficiently, leading to improved second stage processing.
@@
4. Online Evaluation
We introduce our proposed cascade framework to the ICD task aimed at reducing inappropriate content on the short-video platform for our online evaluation.
Model Deployment. The MLLM, fine-tuned during offline training, has been deployed online using A10 GPUs with a 24GB memory capacity. For our online A/B test, thanks to our cascade structure, only up to 5% of the GPUs used for full-traffic MLLM deployment are required.
Online A/B experiment results. We conducted online A/B tests, comparing our cascade serving framework (experiment group) with the online setup serving only the base model (control group). The evaluation metric, the inappropriate content video view rate (ic_vvr), is calculated daily based on human-labeled samples. When a video is identified as inappropriate, a deboosting strategy is applied in the multi-stage recommendation funnel to reduce its visibility, thereby lowering the ic_vvr platform-wide. Over a 10-day experiment with 10% traffic allocated to each group, our cascade framework reduced ic_vvr by 9.9% (statistically significant) with no impact on core metrics.
Post-launch evaluation. Following the launch of the experiment group, the cascade serving system has been running for more than a month. Global metric monitoring shows a substantial and meet-expectation reduction in ic_vvr, confirming the effectiveness of the launch.
@@ 5. Related Work
Additionally, SigLIP (Zhai et al., 2023) introduces a simplified approach by employing a pairwise sigmoid loss for image-text pre-training, providing an alternative to conventional contrastive objectives. BEiT3 (Wang et al., 2022) proposes a general-purpose multimodal foundation model that adopts a multi-way transformer architecture, integrating both masked image modeling and masked language modeling objectives to support a broad array of vision-language tasks.
Audio Modality. OpenAI’s Whisper model (Radford et al., 2023) has demonstrated exceptional speech recognition capabilities by utilizing large-scale weak supervision for speech-to-text tasks. Its training on diverse languages and audio formats enables strong generalization across various recognition challenges, reflecting the growing trend of leveraging massive datasets to enhance deep learning models in audio processing. While video-audio (Liu et al., 2023a; Zhu et al., 2024) and text-audio (Radford et al., 2023) models are increasingly prevalent, pretrained models (Liu et al., 2023b; Sun et al., 2025) that seamlessly integrate all three modalities—video, text, and audio—remain limited, particularly in the context of social media data.
@@
To address this gap, LLaVA-OneVision(Li et al., 2024) has been introduced with the goal of achieving top-tier performance across single-image, multi-image, and video tasks, thereby exhibiting robust generalization and effective transfer of learned features in cross-scenario. By leveraging advanced techniques for cross-modal alignment and representation, LLaVA-OneVision enables flexible adaptation and accurate predictions across different visual contexts
Several other versatile open models have also shown promising potential for multi-scenario applications in this space. Video-LLaMA(Zhang et al., 2023) proposes a multimodal framework that empowers Large Language Models (LLMs) with the capability of understanding both visual and auditory content in the video. VILA(Lin et al., 2024) examined effective pretraining strategies to adapt LLMs for vision-related tasks. InternLMX-Composer-2.5(Zhang et al., 2024) offers advanced capabilities for handling long-context input and output, supporting features like ultra-high resolution image comprehension, detailed video analysis, multi-turn multi-image dialogue, webpage creation, and article composition. Meanwhile, Qwen-VL series(Wang et al., 2024; Bai et al., 2025) introduces the Naive Dynamic Resolution mechanism, enabling the model to dynamically process images of varying resolutions into different quantities of visual tokens. Ovis(Lu et al., 2024) integrates an additional learnable visual embedding table into the visual encoder’s process to structurally align visual and textual embeddings.
@@ 6. Conclusion
Modern video understanding models benefit significantly from a multimodal approach. By leveraging the complementary strengths of these modalities, multi-tower architectures achieve robust and holistic performance across general tasks. To better address the nuanced demands of specialized applications, a unified Multimodal LLM framework can further improve cross-modal interactions and sensitivity to domain-specific cues, such as assessing video quality or policy compliance on short-video platforms, surpassing traditional multi-tower limitations.
Given the substantial GPU demands of Multimodal LLMs, especially in real-time scenarios, we propose a cascade structure for efficient deployment. This two-stage framework begins with a lightweight model that assigns an entropy-based uncertainty score to each video, filtering out low-uncertainty cases before engaging the MLLM for deeper analysis. By prioritizing high-uncertainty inputs, this method significantly reduces computational overhead while maintaining classification accuracy. Empirical results show that this cascade approach not only optimizes GPU use but also enhances classification precision for specialized tasks, marking a substantial advancement in scalable multimodal video classification.
@@ 7. Future Exploration
Future work could extend COEF-VQ into a multi-issue MLLM capable of handling various video quality tasks (e.g., quality classification, content moderation, and genre classification) within a single model. This unified approach would streamline processing, enhance consistency, and reduce the need for multiple task-specific models, optimizing resource and maintenance efforts.
Another key improvement is enhancing audio modality integration. The current late-fused feature concatenation may not fully align text, audio, and visual cues. Future versions could incorporate modality alignment techniques to better capture interactions between these signals, improving the model’s ability to distinguish critical audio cues (e.g., background noise vs. intentional soundtracks) and leading to more precise video quality assessments.